OpenStudy (anonymous):

Which fractions are equivalent to the fraction below? Check all that apply. 6/21 a. 2/22 b. 2/7 c. 2/28 d. 12/42

OpenStudy (anonymous):

3 over 7 because 9/3=3 21/3=7

OpenStudy (anonymous):

i think its more than onee

OpenStudy (anonymous):

wait what 3/7 aint no option

OpenStudy (anonymous):

also d becuase 12/42 /2 2 =d

OpenStudy (anonymous):

2/7